OpenStudy (anonymous):

The long jump record, in feet, at a particular school can be modeled by f(x)=18.1+2.2in(x+1) where x is the number of years since records began to be kept at the school. What is the record for the long jump 14 years after record started being kept? Round your answer to the nearest tenth.

OpenStudy (anonymous):

i am going to guess this is \[f(x)=18.1+2.2\ln(x+1)\]\

OpenStudy (anonymous):

your job is to use a calculator and compute \[f(14)=18.1+2.2\ln(14+1)\]

OpenStudy (anonymous):

@satellite73 wut would this one be can u help me out

OpenStudy (anonymous):

just calculate?

OpenStudy (anonymous):

lol how did you find this??

OpenStudy (anonymous):

google loll

OpenStudy (anonymous):

ok yes, just calculate use wolfram

OpenStudy (anonymous):

i tried usinng that even paid for pro and i think it sucks or am just confused when i put equations in

OpenStudy (anonymous):

type it in as you see it looks like about \(24.1\) rounded to the nearest tenth

OpenStudy (anonymous):

guess y'all had the same on line class, and they didn't bother to change the questions on line summer math = punishment

OpenStudy (anonymous):

yea im getting 24.1 .... n im taking a calculus class

OpenStudy (anonymous):

well this is not really calc, but whatever you got more?

OpenStudy (anonymous):

i know!!!! lol. ya

OpenStudy (anonymous):

lets knock a couple out

OpenStudy (anonymous):

kk ill start a new one

OpenStudy (anonymous):

k